# ModernGBERT: German-only 1B Encoder Model Trained from Scratch

### v1

@@ Abstract
Despite the prominence of decoder-only language models, encoders remain crucial for resource-constrained applications. We introduce ModernGBERT (134M, 1B), a fully transparent family of German encoder models trained from scratch, incorporating architectural innovations from ModernBERT. To evaluate the practical trade-offs of training encoders from scratch, we also present LLäMmlein2Vec (120M, 1B, 7B), a family of encoders derived from German decoder-only models via LLM2Vec. We benchmark all models on natural language understanding, text embedding, and long-context reasoning tasks, enabling a controlled comparison between dedicated encoders and converted decoders. Our results show that ModernGBERT 1B outperforms prior state-of-the-art German encoders as well as encoders adapted via LLM2Vec, with regard to performance and parameter-efficiency. All models, training data, checkpoints and code are publicly available555ModernGBERT and LLäMmlein2Vec, including all code, data and intermediary checkpoints will be published upon acceptance under a “research-only RAIL” license., advancing the German NLP ecosystem with transparent, high-performance encoder models.
1 Introduction
Despite the recent dominance of decoder-only large language models (LLMs), parameter-efficient encoder models remain crucial for language technology, particularly for local deployments such as retrieval-augmented generation (RAG). Their bidirectional attention confers strong understanding capabilities with lower resource requirements, making them attractive for consumer hardware. In the German NLP landscape, GBERTLarge (337M parameters; Chan et al., 2020) remains a popular encoder, performing competitively with much larger German-capable decoder LLMs across tasks (Pfister and Hotho, 2024), despite its modest size and limited training data (163 GB). More recently, ModernBERT (Warner et al., 2024) introduced several architectural improvements for English encoders, including enhanced relative positional embeddings and efficient attention patterns enabling long context processing. Building on this progress and inspired by the success of LLäMmlein (Pfister et al., 2024), a family of German decoder-only LLMs transparently trained on approximately of RedPajamaV2 (Weber et al., 2024) text, we introduce ModernGBERT a family of fully open, high-performance German encoder models with 134M and 1B parameters.
These models provide a foundation to explore the impact of ModernBERT’s architectural innovations on German encoder performance. They also allow us to investigate how parameter scaling influences model quality when trained on large-scale monolingual corpora.
To better assess the practical utility and trade-offs of training encoder models from scratch, we further present LLäMmlein2Vec encoders (120M, 1B, and 7B), derived from decoder-only models using LLM2Vec (BehnamGhader et al., 2024). Since all models are based on the same training datasets, this setup provides a foundation for systematically analyzing the relationship between different architectures and training strategies.
We extensively evaluate these models during and post training via: natural language understanding (SuperGLEBer, Pfister and Hotho, 2024), embedding performance (MTEB; Enevoldsen et al., 2025; Muennighoff et al., 2023; Wehrli et al., 2023), and long-context understanding (Question Answering Needle-in-a-Haystack). Our findings reveal:
ModernGBERT 134M and 1B are highly competitive German encoders, scaling well with size ( tokens), with 1B surpassing the previous SotA GBERTLarge.
Our LLäMmlein2Vec 7B also outperforms GBERTLarge, though dedicated encoders still outperform converted models of similar size.
2 Datasets
2.1 Pre-training Dataset
@@
We pre-trained ModernGBERT on the same data as LLäMmlein decoder models (Pfister et al., 2024), using the open-source RedPajamaV2 dataset Weber et al. (2024).666Common Crawl Foundation Terms of Use This dataset comprises German CommonCrawl snapshots from 2014–2023. As we intend to keep datasets constant between ModernGBERT and LLäMmlein, we follow LLäMmlein’s data pipeline and select the higher quality document-level deduplicated “head” and “middle” partitions, excluding the lower quality “tail” partition. For our 134M model, we only selected the head partition. We used the same processing pipeline as Pfister et al. (2024): First, paragraph-level deduplication using a Bloom filter to remove redundant content like GDPR notices and web boilerplate, improving data diversity. Then, a token-to-word ratio filter to further improve text quality. The final dataset is approximately , consisting of from head and from middle. Using a GBERTLarge tokenizer, this results in about 1.27T tokens.
2.2 Context Extension Dataset
@@
ModernBERT enhances its context capacity from to by finetuning in two phases: on 250B-token subsample of -token sequences from its original pre-training dataset (ext1), followed by a curated 50B-token dataset with mixed sequence lengths, including short and long ones (ext2, up to tokens) Gao et al. (2025).
Following this setup, we proceed to construct our own German context extension datasets for the two phases (Table 1): for the first phase (ext1), we take the same approach and subsample long sequences from our pre-training datasets (resulting in “LONG-Head” from the head partition for our 134M model, and “LONG-Head/Middle” from the head and middle partition for our 1B model). For the second phase (ext2) on a high-quality dataset we coin “HQ” in Table 1, we use the German portion of the Fineweb2 dataset (Penedo et al., 2024)777Open Data Commons Attribution License (ODC-By) v1.0. Aiming for a similar distribution, we first take a randomized sample of Fineweb2, and added a separate sample from Fineweb2, selecting long documents with tokens, splitting them into sequences of tokens (“Fineweb2-long”). Furthermore, we add additional long documents by including the 2023 German Wikipedia888Creative Commons Attribution-ShareAlike 3.0 and the 2022 OpenLegalData dump,999https://de.openlegaldata.io/; database licensed under Open Database License (ODbL v1.0), cases are exempt from copyright law also split to sequences of up to tokens. The entire HQ dataset consists of 14.4B tokens. Table 1 summarizes these three resulting datasets.
Dataset | # tokens | # sequences | median length |
LONG-Head (ext1) | B | ||
LONG-Head/Middle (ext1) | B | ||
HQ (ext2) | B | ||
Fineweb2 | M | ||
OpenLegalData | M | ||
Wikipedia | M | ||
Fineweb2-long | M |
3 Methodology
3.1 ModernGBERT: Scaling SotA Encoders
@@
The ModernGBERT models adapt the ModernBERT architecture and training strategy for German. ModernGBERT 134M matches the base ModernBERT model size (22 layers, 768 hidden units, but 16M fewer parameters due to a smaller vocabulary size), while ModernGBERT 1B consists of 28 layers and a hidden size of . Full architectural details can be found in Table 4.
@@
Both models follow the ModernBERT pre-training recipe: masked language modeling (MLM) with no next-sentence-prediction, a 30% masking rate, and sequences up to tokens ( RoPE theta). For training, we use our German pre-training corpus (Section 2.1): ModernGBERT 1B trains on the head then middle partitions for a total of 1.27T tokens; ModernGBERT 134M is trained only on the head partition (0.47T tokens), as downstream evaluation showed early saturation (Section 5.1). Details of the training procedure are shown in Table 5.
After MLM, we extend context length in two phases, following ModernBERT, raising the RoPE theta to and training on longer sequences. In the first extension phase (ext1), models are trained on the LONG-Head for the 134M model or LONG-Head/Middle for the 1B model. In ext2, both models are trained on the HQ dataset. As we did not intend to develop a novel German tokenizer, we utilized the original BERT-style tokenizer from GBERTLarge (resulting in a -word embedding layer - a multiple of 64 for compute efficiency). While LLäMmlein (Pfister et al., 2024) provides a dedicated German Llama-style tokenizer, our preliminary ablations consistently showed degraded downstream performance. This degradation is consistent with results by Warner et al. (2024), who observed similar behavior with (English) Llama-style tokenizers during the development of the original ModernBERT. We therefore retained the GBERTLarge tokenizer.
@@
Throughout the training checkpoints are saved and evaluated, and all are released publicly to support further research. In addition, inspired by Pythia (Biderman et al., 2023) we provide full training provenance by logging and releasing the order of data points seen during training; thus, all checkpoints can be linked with the exact data points seen up to that checkpoint.
3.2 LLM2Vec: Turning Decoders to Encoders
LLM2Vec BehnamGhader et al. (2024) proposes a method to convert decoder-only LLMs into effective text encoders through the following steps: First, the causal attention mask is replaced with a full attention mask, enabling bidirectional attention across tokens. Second, the model is trained using a masked next token prediction (MNTP) objective. Third, unsupervised contrastive learning (SimCSE) is applied, improving embedding quality by maximizing agreement between differently dropped-out versions of the same input. However, we intend to remain closely aligned with ModernGBERT’s training objectives and not reproduce LLM2Vec results. For this reason, we trained our models exclusively using the MNTP objective, which is most similar to MLM. ModernGBERT performs two context extension phases, each using two datasets per model. Similarly, we train all three LLäMmlein models using the same respective two datasets as employed by ModernGBERT’s context extensions (Section 2.2): the LLäMmlein2Vec 120M model variant follows ModernGBERT 134M (LONG-Head for phase one, HQ for phase two); the LLäMmlein2Vec 1B and 7B models follow ModernGBERT 1B (LONG-Head/Middle for phase one, HQ tokens for phase two). For each model, we apply MNTP training separately on each respective dataset, resulting in two distinct adapter modules—one per phase. We evaluate both individual adapters (ext1 & ext2) as well as a merged model (ext1+2) where both adapters are combined. Notably, the models achieve comparable results even without seeing the full training data as exemplary shown in Table 7 - this is comparable to the observations in Pfister et al. (2024), indicating possibilities of reducing compute in future trainings. However, for consistency and comparability, we report results using the fully trained models throughout the paper101010except for the LLäMmlein2Vec 7B trained on the LONG-Head/Middle model, which we trained on 64 nodes with 4 H200 each for 14 hours, before stopping the training due to compute constraints (Table 6). We also increased the sequence length to and set RoPE theta to , otherwise, we follow the default LLM2Vec parameters (see Table 6 for more details).
4 Evaluation Setup
4.1 SuperGLEBer
@@
We assess our final models using the German SuperGLEBer benchmark Pfister and Hotho (2024), which includes 29 tasks across text classification, sequence tagging, question answering, and sentence similarity. These tasks cover diverse domains such as news, legal texts, and consumer reviews. For each task, models are fine-tuned with QLoRA Dettmers et al. (2023) by default, or LoRA as fallback. In addition to evaluating final checkpoints, we follow LLäMmlein Pfister et al. (2024) and evaluate intermediate checkpoints on the same representative SuperGLEBer subset as selected by Pfister et al.: the classification tasks NLI Conneau et al. (2018), FactClaiming Comments Risch et al. (2021), DB Aspect Wojatzki et al. (2017), and WebCAGe Henrich et al. (2012), the sequence tagging task EuroParl Faruqui and Padó (2010), and the sentence similarity task PAWSX Liang et al. (2020).
4.2 Massive Text Embedding Benchmark
@@
We further evaluate the models on the German subset of the Massive Text Embedding Benchmark MTEB(deu,v1) (Enevoldsen et al., 2025). The specific tasks can be found in Table 8. In addition to text pair classification and semantic textual similarity—already covered by the SuperGLEBer benchmark—MTEB includes clustering Wehrli et al. (2023), as well as reranking and retrieval tasks.
These latter tasks provide a more comprehensive assessment of general-purpose sentence embeddings, focusing on the models’ ability to produce robust semantic representations.
@@
4.3 Long-Context Understanding
Evaluating long-context capabilities in German is hindered by the scarcity of native high-quality datasets, with translations from English often introducing artifacts.
@@ 4.3 Long-Context Understanding
To address this, we construct a Question-Answering Needle-In-a-Haystack (QA-NIAH) evaluation (Ivgi et al., 2023; Hsieh et al., 2024) based on the human-annotated GermanQuAD dataset (Möller et al., 2021). Given a question, the goal is to extract the answer span from a long document. We adapt GermanQuAD to a QA-NIAH setup as follows: for each question–paragraph (“needle”) pair, we sample up to 3 distractor paragraphs and shuffle them with the needle, forming a “haystack” document of up to tokens. The answer always appears only in the needle paragraph. For evaluation, we increase distractors to up to 20, yielding documents up to tokens. This yields training and test question–haystack pairs. We fine-tune models on the QA-NIAH training set using QLoRA, and evaluate on the test set, following the SuperGLEBer benchmark procedure. Training uses sequences up to tokens; evaluation uses up to tokens, assessing generalization to longer contexts.
5 Evaluation Results
5.1 Intermediate Model Evaluation
To track pre-training progress, we evaluated intermediate checkpoints on six representative SuperGLEBer tasks, following Pfister et al. (2024) (see Section 4.1). All checkpoints will be released for future analysis. Figure 3 shows that ModernGBERT 1B’s average performance steadily improves throughout training, while ModernGBERT 134M quickly saturates.
To quantify these trends, we evaluated the full SuperGLEBer suite on four checkpoints for ModernGBERT 1B and three for 134M (comparison done using Wilcoxon signed-rank tests): ModernGBERT 134M plateaued after 72B tokens (15% of data), with no further significant improvements. In contrast, ModernGBERT 1B showed significant gains over the same dataset portion (), and additional gains during training on the middle partition (). Performance then plateaus after B tokens (67% of entire pre-training dataset), with the SuperGLEBer score increasing only slightly from 0.777 to 0.791 despite 406B more tokens.
Analyzing on the six selected subtasks that were run on every checkpoint (Section 5.1), for the 134M variant, only PAWSX showed a significant positive Spearman rank correlation between training token count and performance (), while the others did not. For 1B, all but EuroParl showed significant positive correlations (). In particular, even though the aggregate score remains largely stable in the final third of the pre-training, on complex tasks such as NLI and PAWSX, we still see slight improvements with increased training (Figure 2).

### v2

# New Encoders for German Trained from Scratch: Comparing ModernGBERT with Converted LLM2Vec Models

@@ Abstract
Encoders remain essential for efficient German NLP and NLU scenarios despite the rise of decoder-only LLMs.
This work studies two routes to high-quality German encoders under identical data and training constraints:
a) training from scratch and
b) converting decoders via LLM2Vec.
We introduce two resources:
ModernGBERT (134M, 1B), fully transparent German encoders in the ModernBERT style, and
LLäMmlein2Vec (120M, 1B, 7B), decoder-to-encoder conversions trained with masked next-token prediction,
both undergoing a context extension to tokens.
Across SuperGLEBer, ModernGBERT 1B sets a new state of the art (avg 0.808), surpassing GBERTLarge (+4%) and the seven-times larger converted 7B model (0.787).
On German MTEB after supervised fine-tuning, ModernGBERT 1B (0.551) approaches the converted 7B model (0.557).
We release all models, checkpoints, datasets, and full training records, and introduce an encoder-adapted QA-NIAH evaluation.
All in all, our results provide actionable guidance:
when parameter efficiency and latency matter, from-scratch encoders dominate. When a pre-trained decoder exists and compute is a limited, conversion offers an effective alternative.111ModernGBERT and LLäMmlein2Vec, including all code, data and intermediary checkpoints will be published upon acceptance under a research-only RAIL license.
Keywords: Language Resources, Encoder-only Models, German, Model Conversion, Long Context, Benchmarking
New Encoders for German Trained from Scratch:
Comparing ModernGBERT with Converted LLM2Vec Models
| Julia Wunderle1,∗, Anton Ehrmanntraut2,∗, Jan Pfister1 |
| Fotis Jannidis2,†, Andreas Hotho1,† |
| 1Data Science 2Computer Philology and History of Contemporary German Literature |
| CAIDAS – Center for Artificial Intelligence and Data Science |
| JMU – Julius-Maximilians-Universität Würzburg |
| {lastname}@informatik.uni-wuerzburg.de, {firstname.lastname}@uni-wuerzburg.de |
| Equal contribution |
Abstract content
1. Introduction
Encoders remain central to efficient (German) NLP and NLU where latency, memory, and cost dominate Pfister and Hotho (2024), still accounting for most of the downloads from e.g. Hugging Face Bourdois (2025). Especially for tasks like sentence similarity (i.e. important for RAG), mask-filling and text/token classification, encoder models are still the architecture of choice Bourdois (2025). Yet, the German NLP landscape lacks high-quality, modern, transparent encoders with reproducible training provenance and long-context support. This paper contributes both resources and a controlled study addressing two practical routes to encoders under identical data and training constraints: training from scratch versus converting decoders.
We introduce to complementary model families: ModernGBERT (134M, 1B), ModernBERT-style German encoders trained from scratch, and LLäMmlein2Vec (120M, 1B, 7B), encoders derived from German decoder-only models via LLM2Vec. ModernBERT (Warner et al., 2024) introduced several architectural improvements for English encoders, including improved relative positional embeddings and efficient attention patterns that allow long context processing. We adapt this design to German, providing strong encoder baselines built entirely from scratch.
In parallel, to assess the practical utility and trade-offs of training encoder models from scratch, we converted the decoder-only model family LLäMmlein into encoders using LLM2Vec (BehnamGhader et al., 2024). Specifically, to align closely with the encoder training objective, we limit the procedure to the first two LLM2Vec steps: 1. enabling a bidirectional attention mask, and 2. masked next-token prediction (MNTP). Since all models share the same training datasets, this setup provides a unique foundation for systematically analyzing the relationship between different architectures and training strategies. We extensively evaluate and compare these models during and post training via: natural language understanding (SuperGLEBer, Pfister and Hotho, 2024), embedding performance (MTEB; Enevoldsen et al., 2025; Muennighoff et al., 2023; Wehrli et al., 2023), long-context understanding (new Question Answering Needle-in-a-Haystack (QA-NIAH)) and an efficiency suite reflecting variable-length inference. Our key contributions are:
We introduce a ModernGBERT family, which achieves new state-of-the-art performance on SuperGLEBer and the German MTEB.
To enable comparisons between training strategies, we also introduce a decoder-turned-encoder LLäMmlein2Vec family, based on the same training dataset.
We find that dedicated encoders trained from scratch consistently outperform converted decoders of similar size.
Our newly introduced QA-NIAH benchmark confirms strong long-context understanding of our models, validating the effectiveness for extended input sequences.
To enable further research we release all (intermediary) resources including all model checkpoints, data point tracking, code and data.
2. Datasets
2.1. Pre-Training Dataset
@@
We pre-trained ModernGBERT on the same data as LLäMmlein decoder models (Pfister et al., 2024), using the open-source RedPajamaV2 dataset Weber et al. (2024) . This dataset comprises German CommonCrawl snapshots from 2014 to 2023. As we intend to keep datasets constant between ModernGBERT and LLäMmlein, we follow LLäMmlein’s data pipeline and select the higher quality document-level deduplicated “head” and “middle” partitions, excluding the lower quality “tail” partition. For our 134M model, we only selected the head partition. Our processing pipeline mirrors Pfister et al. (2024): we first perform paragraph-level deduplication using a Bloom filter to remove redundant content such as GDPR notices and web boilerplate, enhancing data diversity. Afterward, we apply a token-to-word ratio filter to further improve text quality. The final dataset is about , corresponding to approximately 1.27T tokens using a GBERTLarge tokenizer.
2.2. Context Extension Datasets
@@
ModernBERT enhances its context capacity from to by fine-tuning in two phases: on an 250B-token subset of -token sequences from the original pre-training data (ext1), followed by 50B-token high-quality dataset with mixed sequence lengths (ext2) Gao et al. (2025).
Following this setup, we construct two German datasets. For ext1, we analogously subsample long sequences from our pre-training dataset, resulting in “LONG-Head” from the head partition for our 134M model, and “LONG-Head/Middle” from the head and middle partition for our 1B model.
For the high-quality dataset used for ext2, we selected the German portion of Fineweb2 as basis (Penedo et al., 2024) . To match the original distribution, we take a randomized sample of Fineweb2, and add a separate Fineweb2 sample, selecting long documents with tokens, splitting them into sequences of about tokens (“Fineweb2-long”). Additional long documents are drawn from the 2023 German Wikipedia and the 2022 OpenLegalData dump, . The resulting dataset contains 14.4B tokens, summarized in Table˜1.
| Dataset | # Tokens | # Sequences | Median Length | |
ext1 |
LONG-Head | B | ||
| LONG-Head/Middle | B | |||
ext2 |
High Quality | B | ||
M | |||
M | |||
M | |||
M |
3. Methodology 111All training settings and parameters will be detailed in the camera-ready version.
3.1. ModernGBERT
@@
We adapt the ModernBERT architecture and training strategy for German. While ModernGBERT 134M matches the base ModernBERT model size (22 layers, 768 hidden units, but 16M fewer parameters due to a smaller vocabulary size), we create ModernGBERT 1B with 28 layers and a hidden size of .
Pre-/Post-Training Strategy
@@
Both models are pre-trained using masked language modeling (MLM) with no next-sentence-prediction, a 30% masking rate, and sequences up to tokens ( RoPE theta). ModernGBERT 134M is trained on the head partition (0.47T tokens), as downstream evaluation indicated early saturation, while ModernGBERT 1B is trained on both the head and middle partitions of our pre-training corpus (Section˜2.1), totaling 1.27T tokens.
After standard MLM pre-training, we proceed with the two context extension phases: During ext1, the RoPE theta is raised to , and models are trained on the LONG-Head (134M) or LONG-Head/Middle (1B) datasets (see Section˜2.2). In the ext2 phase, both models are trained on the High Quality dataset.
Tokenization
For tokenization, we use the original BERT-style tokenizer from GBERTLarge, resulting in a -word embedding layer. While LLäMmlein (Pfister et al., 2024) provides a dedicated German BPE tokenizer, our preliminary ablations using this tokenizer consistently showed degraded downstream performance. Consequently, we retained the GBERTLarge tokenizer.
Training Progress Tracking
@@
Throughout the training we save, evaluate and release all checkpoints to support further research. Inspired by Pythia (Biderman et al., 2023) we provide full training transparency by logging and releasing the order of data points seen during training; thus, all checkpoints can be linked with the exact data points seen up to that checkpoint.
3.2. LLM2Vec: Turning Decoders to Encoders
BehnamGhader et al. (2024) proposes a method to convert decoder-only LLMs into text encoders through the following steps: First, the causal attention mask is replaced with a full attention mask, enabling bidirectional attention across tokens. Second, the model is trained using a masked next token prediction (MNTP) objective. Third, regular LLM2Vec includes a unsupervised contrastive learning (SimCSE) step, improving embedding quality by maximizing agreement between differently dropped-out versions of the same input. However, we intend to remain as close as possible to the training objectives of ModernGBERT to allow fair and direct comparisons. Therefore, we limit the procedure to the first two steps of LLM2Vec, specifically employing the MNTP objective, which is most closely aligned with MLM.
Pre-/Post-Training Strategy
We train all three LLäMmlein models (120M, 1B, 7B) using the same two context extension datasets as employed by ModernGBERT’s context extensions (ext1 and ext2) (Section˜2.2). Specifically, the 120M model mirrors ModernGBERT 134M, being trained on the LONG-Head (ext1) and High Quality (ext2) datasets, while the 1B and 7B models follow ModernGBERT 1B, using LONG-Head/Middle (ext1) and the High Quality dataset (ext2). Notably, MNTP training is applied separately to each dataset, resulting in two adapter modules for each model. Additionally, all adapters were trained on the full corresponding datasets333except for the LLäMmlein2Vec 7B trained on the LONG-Head/Middle model, which we trained on 64 nodes with 4 H200 each for 14 hours, before stopping the training due to compute constraints., although models were able to achieve comparable results during earlier training stages. This indicates, that compute time could have been drastically reduced.
Further aligning with ModernGBERT we combined MNTP with context extension — a setup that, to our knowledge, has not yet been widely explored. Therefore, we also extended the model sequence length to tokens and increased RoPE theta to . We evaluate both individual adapters (ext1 & ext2) as well as a merged model (ext1+2) that combines both
4. Evaluation Setup
4.1. SuperGLEBer
@@
We assess our final models using the German SuperGLEBer benchmark Pfister and Hotho (2024), which includes 29 tasks across text classification, sequence tagging, question answering, and sentence similarity. These tasks cover diverse domains such as news, legal texts, and consumer reviews. For each task, models are fine-tuned with QLoRA Dettmers et al. (2023) by default, or LoRA as fallback. In addition to evaluating final checkpoints, we follow LLäMmlein Pfister et al. (2024) and evaluate intermediate checkpoints of ModernGBERT as well as LLäMmlein2Vec on the same representative SuperGLEBer subset consisting of: the classification tasks NLI Conneau et al. (2018), FactClaiming Comments Risch et al. (2021), DB Aspect Wojatzki et al. (2017), and WebCAGe Henrich et al. (2012), the sequence tagging task EuroParl Faruqui and Padó (2010), and the sentence similarity task PAWSX Liang et al. (2020).
4.2. Massive Text Embedding Benchmark
@@
We further evaluate the models on the German subset of the Massive Text Embedding Benchmark MTEB(deu,v1) (Enevoldsen et al., 2025), comprising 19 tasks. In addition to text pair classification and semantic textual similarity, which are already covered by the SuperGLEBer benchmark, MTEB includes clustering Wehrli et al. (2023), as well as reranking and retrieval tasks. These latter tasks provide a more comprehensive assessment of general-purpose sentence embeddings, focusing on the models’ ability to produce robust semantic representations.
@@
4.3. Long-Context Understanding
@@
Evaluating long-context capabilities in German is hindered by the scarcity of native high-quality datasets, with translations from English often introducing artifacts. To address this, we construct a new Question-Answering Needle-In-a-Haystack (QA-NIAH) evaluation (Ivgi et al., 2023; Hsieh et al., 2024) based on the human-annotated GermanQuAD dataset (Möller et al., 2021). To our knowledge, this is the first QA-NIAH evaluation specifically adapted for encoder models. The goal in QA-NIAH is to extract an answer span from a long document. We adapt GermanQuAD to a QA-NIAH setup as follows: for each question-paragraph (“needle”) pair, we sample up to 3 distractor paragraphs and shuffle them with the needle, forming a “haystack” document of up to tokens. The answer always appears only in the needle paragraph. For evaluation, up to 20 distractors are included to test generalization to longer contexts, yielding documents up to tokens. This results in training and test question–haystack pairs. Models were again fine-tuned using QLoRA falling back to LoRA, where quantization is not supported. We will release the QA-NIAH generation code and splits to facilitate reuse and extension.
5. Evaluation Results
The tables in this paper present only a summary of our results. Extensive benchmarking results, analyses, and all hyperparameters of all models will be provided in the camera-ready version.
5.1. Intermediate Model Evaluation
ModernGBERT Training
To track the pre-training progress, we evaluated intermediate checkpoints on the SuperGLEBer benchmark (see Section˜4.1). Figure˜2 exemplary shows the results for the ModernGBERT 1B model in the top row. To quantify trends, we first assessed the full SuperGLEBer suite on five 1B and respective three 134M equally spaced ModernGBERT checkpoints and performed Wilcoxon signed-rank tests. The 134M model plateaued after 72B tokens (15% of data), with no further significant gains. In contrast, the 1B model achieved significant improvements after the same amount of data (p 0.0001) and on the middle partition (p 0.00052), before plateauing after 864B tokens (67% of data), with only minor gains thereafter (0.777 0.791)
To gain more insights, we tracked six representative SuperGLEBer subtasks at each checkpoint. For the 134M model, only PAWSX showed a positive Spearman rank correlation with training duration (r = 0.655; p 0.003), whereas for 1B, all tasks except EuroParl improved significantly (r 0.57; p < 0.00014), and complex tasks like PAWSX continued to show modest gains even after the aggregate score stabilized (gray in Figure˜2).